OpenStudy (anonymous):

are 3x + 33/ x+ 11 and 3 equivalent? Using complete sentences explain why or why not.

OpenStudy (diyadiya):

I can show you how

OpenStudy (anonymous):

Thanks!

OpenStudy (diyadiya):

\[\frac{3x+33}{x+11}= \frac{3(x+11)}{(x+11)}=\frac{3\cancel{(x+11)}}{\cancel{(x+11)}}=3\]

OpenStudy (diyadiya):

When we factor out 3 from 3x+33 we get 3(x+11) understood?

OpenStudy (anonymous):

yes

OpenStudy (diyadiya):

3(x+11) =3*x + 3*11 =3x + 33 ok ?

OpenStudy (anonymous):

ok

OpenStudy (diyadiya):

:)

OpenStudy (diyadiya):

So they are equivalent

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

so how can i explain that in words

OpenStudy (diyadiya):

Sorry My explainations aren't good enough

OpenStudy (anonymous):

its perfectly fine i got it can you help me with something else please

OpenStudy (anonymous):

Using complete sentences, explain each step in simplifying the ratio x^2 + 8x -------------- x^2 + 10x + 16

OpenStudy (anonymous):

and provide the simplified form.

OpenStudy (diyadiya):

You have to factorise both equations x^2 + 8x & x^2 + 10x + 16

OpenStudy (diyadiya):

How do you factorise x^2 + 10x + 16 ?

OpenStudy (anonymous):

(x+8)(x+2)

OpenStudy (diyadiya):

Yes Exactly!!

OpenStudy (diyadiya):

Now x^2 + 8x

OpenStudy (anonymous):

x(x+8)

OpenStudy (diyadiya):

Yes Now put it in a fraction ....& cancel out (x+8) since its common

OpenStudy (anonymous):

alright then what do you get

OpenStudy (diyadiya):

x(x+8) ---------- (x+8)(x+2) what do you get after cancelling (x+8) ?

OpenStudy (anonymous):

um not sure

OpenStudy (diyadiya):

Try ,its ok if you're wrong

OpenStudy (anonymous):

i gotta go.. i gotta go get dinner.. its getting late

OpenStudy (diyadiya):

Oh ok

OpenStudy (diyadiya):

cya later then :)